OpenStudy (anonymous):

Two pipes A and B can fill a cistern in 371/2 minutes and 45 minutes respectively. Both pipes are opened. The cistern will be filled in just half an hour, if the B is turned off after

OpenStudy (anonymous):

37 and 1/2 hours

OpenStudy (anonymous):

Look at this formula : To find the h they work togehether : 1/h1+1/h2+....1/hn=1/h____>h/1 OK ?!

OpenStudy (anonymous):

k

OpenStudy (anonymous):

75/2+1/45=1/30

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

11/225=30

OpenStudy (anonymous):

I realy dont understand your numbers :( Can u write them for example : h1=n h2=r OK ?!

OpenStudy (anonymous):

h1=2/75,h2=1/45and h=1/30

OpenStudy (anonymous):

look : h1=275/100____>1/h1=1/175/100=100/275 . OK ?!

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."